Marginal Productivity and Marginal Cost

Crepes

Work hours

Mworkers Or the change in Work hours

MC (at wage of $20/hour), inputs at $5

MPL  amount that one worker would produce

1

0.05

2

0.2

3

0.45

4

0.8

5

1.25

6

1.8

7

2.45

8

3.2

9

4.05

10

5

  1. In the short-run, how can you increase production?  
  2. Fill in the table.  After the first dinner, what happens to the time needed for each additional crepe? Why? What happens to the marginal product of labor?  What happens to marginal costs?
Answer #1

Before filling in the table, let us go through all the formulas we are going to use.

Mworkers = Change in work hours

MC = wage rate * Mworkers +inputs

MPL =Wage rate / MC

Crepes work hours Mworkers

MC(at wage of

$20/hr)

Inputs at $5

MPL amount that

One worker

Would produce

1 0.05 0.05 $(20*0.05+5)=6 (20/6)=3.33
2 0.2 (0.2-0.05)=0.15 $(20*0.15+5)=8 2.5
3 0.45 0.25 10 2
4 0.8 0.35 11.4 1.75
5 1.25 0.45 14 1.43
6 1.8 0.55 16 1.25
7 2.45 0.65 18 1.11
8 3.2 0.75 20 1
9 4.05 0.85 22 0.91
10 5 0.95 24 0.83

In the short run we can increase production by increasing the number of Workers.